# Detecting ionized bubbles around luminous sources during the reionization era using HI 21-cm signal

### v2

@@ Abstract
Measuring the properties of the intergalactic medium (IGM) and sources during the Epoch of Reionization (EoR) is of immense importance. We explore the prospects of probing the IGM and sources through redshifted 21-cm observations of individual ionized bubbles surrounding known luminous sources during the EoR. Accordingly, we simulate HI 21-cm maps, foreground contaminants, and system noise which are specific to the uGMRT and SKA1-Low observations. Following the subtraction of the foreground from the total visibility, we employ a visibility-based matched filter technique to optimally combine the desired HI 21-cm signal while minimizing the system noise. Our analysis suggests that these ionized bubbles can be detected with more than $5\sigma$ significance using approximately $\sim 2000$ and $\sim 3000$ hours of observation time with the uGMRT at redshift $7.1$ and $8.3$, respectively, when the mean neutral hydrogen fraction outside the targeted bubble is $\bm{\sim 0.9}$. The SKA1-Low should be able to detect these bubbles with more than $8\sigma$ significance using only $\sim 100$ hrs of observations. The total observing time increases both for the uGMRT and SKA1-Low when the mean neutral hydrogen fraction outside the targeted bubble decreases. Further, we investigate the impact of foreground subtraction on the detectability and find the signal-to-noise ratio decreases when smaller bandwidth is used. More importantly, we show that the matched filtering method can measure ionized bubble radius and constrain HI-neutral fraction reasonably well, providing deeper insights into the source properties and the intergalactic medium.
@@ 1 Introduction
The outline of the paper is as follows: Section 2 provides a brief overview of the basics of the HI 21-cm signal around ionized bubbles. In section 3, we explain our methods to simulate the HI 21-cm signal around ionized bubbles, foreground contaminants, the generation of visibilities at baselines specific to a particular experiment, and system noise. Section 4 discusses foreground subtraction from the total visibilities and the application of the matched filter technique. Our results are presented in Section 5, while Section 6 presents a summary of the work undertaken in the paper alongside a discussion. Throughout our analysis we use the cosmological parameters $h=0.7$, $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$, $\bm{\sigma_{8}=0.83}$ and $\bm{n_{\rm s}=0.96}$ consistent with the WMAP measurements [45].
@@ 3.1 HI 21-cm signal around ionized bubbles
Here, we briefly describe simulated ionized regions surrounding a very bright quasar at $z=7.1$, and around a collection of galaxies at $z=8.3$ during the EoR. First, we simulated the dark-matter distribution using the $N$-body code [^2] https://github.com/rajeshmondal18/N-body developed by [51].  We simulated this within a cubic box of the comoving size of $\sim 527$ Mpc with $2048^{3}$ grids and the same number of particles. The minimum halo mass resolved in this simulation is $\bm{\sim 6.26\times 10^{9}M_{\odot}}$, corresponding to a minimum of $\bm{10}$ dark matter particles, each with a mass of $\bm{\sim 6.26\times 10^{8}M_{\odot}}$. Then we use a Friends-of-Friends  [^3] https://github.com/rajeshmondal18/FoF-Halo-finder (FoF) [52] halo finder to identify the collapsed halos that host the galaxies/bright quasars. These galaxies/quasars serve as the sources of ionizing photons during the EoR. Subsequently, the resolution of simulated cubes is degraded to $512^{3}$ grids. This resolution resembles the corresponding angular resolution of the uGMRT and SKA1-Low and has the added benefit of reduced computational requirements for simulation. We then employ a semi-numerical prescription [53, 54, 55] on these outputs to simulate ionization maps and the HI 21-cm differential brightness temperature maps.
@@
In the above two cases, the targeted HII bubbles are isolated and nearly spherical. This geometry arises because the ionizing radiation from other sources is negligible, and the IGM outside the targeted bubbles remains highly neutral. While these scenarios may be suitable at the onset of reionization, they are likely unrealistic for the later stages of the process. At later stages of reionization, ionized bubbles around very bright sources are surrounded by neighboring ionized regions, leading to an aspherical shape for the targeted bubble. The third scenario accounts for this complexity. The bottom panel of Fig. 1 displays a 2D map of the differential brightness temperature ($\bm{\delta T_{b}}$) at redshift $\bm{z=7.1}$, where the central targeted HII bubble appears highly aspherical. The quasar at the center is the same as the first case. Surrounding it there are numerous other HII regions, unlike the first case. This results in a much lower mean neutral fraction, which is $\bm{0.52}$ in this scenario.
@@
Figure 1: The top left and bottom panels show differential brightness temperature maps of the HI 21-cm signal around an ionized bubble of comoving size $23.5$ Mpc at $z=7.1$ corresponding to observing frequency of $175$ MHz. The top right panel shows a similar map around an ionized bubble of comoving size $27.8$ Mpc at $z=8.3$ corresponding to observing frequency of $153$ MHz. The mass-averaged neutral hydrogen fraction resulting from the galaxies outside the central HII region is $\bm{0.88}$, $\bm{0.94}$, and, $\bm{0.52}$ for the top-left, top-right and bottom panels respectively. The angular size of simulated maps is $\sim 3.3^{\circ}\times 3.3^{\circ}$ with $\sim 23^{\prime\prime}$ resolution.
@@ 3.2.1 Extragalactic radio point sources
One of the most significant foreground contaminants in HI 21-cm signal observation arises from the extragalactic radio point sources, such as Active Galactic Nuclei (AGN), quasars, and radio galaxies. At lower frequencies ($\bm{\sim 150}$ MHz), these sources dominate the sky at smaller angular scales $\bm{\lesssim 4^{\circ}}$ [30]. Therefore, in order to study the HII regions at frequencies of $\bm{150}$ and $\bm{175}$ MHz, it is essential to model the point sources efficiently.
In our simulations, we model the population of extragalactic radio point sources using the following $\bm{150}$ MHz differential source count model from [58]:
$\bm{\frac{dN}{dS}=\frac{10^{3.75}}{\rm Jy.Sr}\left(\frac{S}{1\,\rm Jy}\right)^{-1.6}},$ (3.3)
where $\bm{S}$ is the flux density of the point sources. We consider the point sources to be in the flux range between $\bm{0.1}$ mJy to $\bm{1}$ Jy and they are randomly distributed across an angular size of $\bm{\sim 3.3^{\circ}\times 3.3^{\circ}}$. Based on these parameters, we obtained $\bm{7775}$ point sources from the simulation. We model the spectral energy distribution of the point sources using a simple power law ($\bm{S\propto\nu^{-\alpha}}$), where the spectral index $\bm{\alpha}$ for each source is randomly assigned within the range of $\bm{0.7}$ to $\bm{0.8}$. The right panel of Figure 2 shows a 2D map of the point sources randomly distributed across an angular area of approximately $\bm{3.3^{\circ}\times 3.3^{\circ}}$, with an angular resolution of $\bm{\sim 23^{\prime\prime}}$.
@@ 3.4 Simulating visibilities
We generate the mock visibility data of the HI 21-cm signal and the foregrounds (DGSE and extragalactic point sources) at all simulated baselines at $256$ frequency channels, centered around frequencies $\nu=175$ MHz and $153$ MHz, using the Discrete Fourier Transformation (DFT) of the total simulated maps. We keep the baseline distribution the same in all frequency channels for simplicity.
@@ 3.5 System noise
Our mock simulation considers $8$ hrs of observations with $16$s and $320$s integration time for the uGMRT and SKA1-Low respectively. This produces a total of $783000$ and $11773440$ baselines for the uGMRT and SKA1-Low respectively. Lower integration time will produce a higher number of baselines for the SKA1-Low and hence will increase the requirement of computational resources which is beyond our reach. The system noise contribution for $8$ hrs of observations will be too high to detect the HI 21-cm signal around the individual ionized bubbles. Therefore, we present our results for $2048$ and $3072$ hrs of uGMRT observations at redshifts $7.1$ and $8.3$, respectively, assuming a total of $256$ and $384$ nights of observations, with each night lasting $8$ hours. Further, for the third scenario, we present the results at redshift $\bm{7.1}$ for $\bm{5120}$ hours, assuming $\bm{640}$ nights of observations. The increased observing time for the third scenario is due to the lower mean neutral hydrogen fraction outside the targeted bubble. Assuming all nights are the same, we reduce the system noise corresponding to an $8$ hrs observation by a factor of $16$, $19.6$, and $\bm{25.29}$ for the three scenarios, respectively. The resulting mock signal becomes equivalent to $2048$, $3072$, and $\bm{5120}$ hours of uGMRT observations. Similarly, for the SKA1-Low we present our results for $96$ hrs of observation for the first two cases at redshift $7.1$ and $8.3$ (neutral hydrogen fraction $\bm{\sim 0.9}$) and for $\bm{200}$ hrs of observation for the third case at redshift $\bm{7.1}$ (neutral hydrogen fraction $\bm{\sim 0.5}$). Eventually, we obtain mock visibility data with all the signal contributions (HI 21-cm signal, galactic foreground, extragalactic point sources, and system noise contribution) typical of the uGMRT and SKA1-Low observations.
@@
Figure 6 shows the total mock visibility comprising of HI signal, diffused galactic synchrotron emission, radiation from extra-galactic radio point sources, and system noise as a function of frequency for two baselines $\bm{U=34\lambda}$ (left panel) and $\bm{U=230\lambda}$ (right panel) for the uGMRT at $\bm{175}$ MHz. We note that Figure 6 shows contribution both from the extragalactic point sources, galactic diffuse synchrotron radiation along with other components unlike Fig. 5 which shows galactic diffuse synchrotron emission only as the foreground component. The radio point sources are modeled using a simplistic power-law form as discussed in subsection 3.2.1. The green lines indicate visibilities due to point sources. We see that at a small baseline (left panel) the diffuse galactic synchrotron emission dominates over the point sources; at a large baseline (right panel), the point sources dominate over the diffuse galactic synchrotron emission. The total mock visibilities are similar for the second scenario at $\bm{z=8.3}$ for both experiments, except for the fact that the system noise and foreground contributions are higher. Therefore, we do not show them here explicitly. The subsequent section discusses the subtraction of the foreground contribution.
@@
Figure 6: Left panel shows the total mock visibilities (HI 21-cm signal + DGSE + radio point sources + system noise) as a function of observing frequency for the uGMRT at baseline $\bm{U=34\lambda}$ and the right panel shows the same at baseline $\bm{U=230\lambda}$ at $\bm{175}$ MHz observing frequency. Red dashed lines show diffuse galactic foreground contamination and the green lines show the point source contribution separately.
@@ 4.1 Foreground subtraction
Figure 7 presents the residual visibility at 175 MHz after the foreground contribution is subtracted from the total mock visibility for the uGMRT (top panels) and the SKA1-Low (bottom panels). Here, we have only considered the diffuse galactic synchrotron emission as the foreground contamination. However, we get a very similar residual signal when we add the radio point source contribution to the foreground contribution. The residual signal contains the HI 21-cm signal, the system noise, and the un-subtracted foreground, if any. To assess the reliability of foreground subtraction, we compare the combined input HI 21-cm signal and system noise with the extracted signal for two different baselines ${U=34\lambda}$ (top left panel) and ${230\lambda}$ (top right panel), as depicted in Figure 8. In the bottom panels of Figure 8, we show a similar comparison for SKA1-Low for two different baselines ${U=11\lambda}$ (bottom left panel) and ${U=254\lambda}$ (bottom right panel). The plots show that the residual visibility closely matches the input visibility data, with only minor deviations. These slight deviations are possibly due to some residual foreground contribution and subtraction of the HI 21-cm signal and the system noise during the foreground subtraction. Consequently, we can conclude that we have mostly subtracted the foreground data from the mock visibility data, making it ready for further analysis, such as matched filtering. We do not present similar results for the second scenario at $z=8.3$ as it is similar to the first scenario.
Figure 8: These plots compare the recovered visibility (red dashed lines) with the original visibility consisting of the HI 21-cm signal and noise contributions (blue lines) for the uGMRT (top panels) and SKA1-Low (bottom panels) at same baselines and observing frequency considered in Figure 5. Here visibilities are plotted for every fourth frequency channel for visual clarity.
@@ 5 Results
Here, we present our findings for the third case, where the targeted ionized bubble is surrounded by other ionized regions which result in a lower neutral hydrogen fraction of $\bm{x_{\rm HI}=0.52}$. Further, the targeted bubble at the center also became aspherical. A lower neutral hydrogen fraction means that the background HI 21-cm signal is weaker, which results in a lower SNR value. We use the same matched filtering technique as described earlier but increased the observation time to $\bm{5120}$ hours for the uGMRT and $\bm{200}$ hours for the SKA1-Low, considering the lower SNR and increased difficulty in detecting ionized regions at a lower neutral fraction. Figure 15 shows the SNR as a function of filter size $\bm{R_{\rm f}}$ for $\bm{5120}$ hours of uGMRT observation (left panel) and $\bm{200}$ hours of SKA1-Low observations (right panel) at $\bm{z=7.1}$. As in previous cases, each curve represents a different noise realization. We see that for the uGMRT, the SNR peaks at a filter size $\bm{\sim 25}$ Mpc, and for the SKA1-Low, it peaks at $\bm{\sim 23}$ Mpc, with minor deviations in different noise realizations.
@@
Figure 15: This figure shows the signal-to-noise ratio for $\bm{10}$ independent noise realizations corresponding to $\bm{5120}$ hours of observation time for the uGMRT (left panel) and $\bm{200}$ hours of SKA1-Low observations (right panel) at redshift $\bm{7.1}$.
@@
Further, we see that peak SNR values in all ten realizations are $\bm{\gtrsim 3}$ for the uGMRT and $\bm{\sim 11}$ for the SKA1-Low. Like the previous cases, we assume a total bandwidth of $\bm{16}$ MHz and $\bm{6.25}$ MHz while subtracting the foreground for the uGMRT and SKA1-Low, respectively. We see that it is possible to reliably detect individual ionized bubbles in the HI 21-cm maps using the uGMRT with $\bm{5120}$ hours of observation time and the SKA1-Low with only $\bm{\sim 200}$ hours of observations at $\bm{z=7.1}$ with significantly much lower neutral hydrogen fraction value.
@@
Figure 16: The left panel shows the mean signal-to-noise (SNR) ratio of $\bm{100}$ independent noise realizations as a function of filter size for $\bm{5120}$ hours of the uGMRT observations at redshift $\bm{7.1}$. The right panel shows the SNR of $\bm{20}$ independent noise realizations for $\bm{200}$ hours of the SKA1-Low observations. The shaded region shows $\bm{1\sigma}$ uncertainty estimated from independent realizations.
[…]
@@
The left panel of Figure 16 shows the mean SNR along with the $\bm{1-\sigma}$ spread for $\bm{100}$ noise realizations of the uGMRT, while the right panel shows the same for the SKA1-Low with $\bm{20}$ realizations at $\bm{z=7.1}$ for $\bm{x_{\rm HI}=0.52}$. Further, the peak location varies between $\bm{16}$ and $\bm{32}$ Mpc for the uGMRT for different realizations (Figure 15), which can be seen in Figure 17, that shows a histogram of peak locations for $\bm{100}$ realizations for the uGMRT. The peak value of the SNR for the uGMRT also changes in the range of $\bm{\sim 3-5}$ and $\bm{10.5-12.5}$ for the SKA1-Low.

### v3

@@ Abstract
Measuring the properties of the intergalactic medium (IGM) and sources during the Epoch of Reionization (EoR) is of immense importance. We explore the prospects of probing the IGM and sources through redshifted 21-cm observations of individual ionized bubbles surrounding known luminous sources during the EoR. Accordingly, we simulate HI 21-cm maps, foreground contaminants, and system noise which are specific to the uGMRT and SKA1-Low observations. Following the subtraction of the foreground from the total visibility, we employ a visibility-based matched filter technique to optimally combine the desired HI 21-cm signal while minimizing the system noise. Our analysis suggests that these ionized bubbles can be detected with more than $5\sigma$ significance using approximately $\sim 2000$ and $\sim 3000$ hours of observation time with the uGMRT at redshift $7.1$ and $8.3$, respectively, when the mean neutral hydrogen fraction outside the targeted bubble is ${\sim 0.9}$. The SKA1-Low should be able to detect these bubbles with more than $8\sigma$ significance using only $\sim 100$ hrs of observations. The total observing time increases both for the uGMRT and SKA1-Low when the mean neutral hydrogen fraction outside the targeted bubble decreases. Further, we investigate the impact of foreground subtraction on the detectability and find the signal-to-noise ratio decreases when smaller bandwidth is used. More importantly, we show that the matched filtering method can measure ionized bubble radius and constrain HI-neutral fraction reasonably well, providing deeper insights into the source properties and the intergalactic medium.
@@ 1 Introduction
The outline of the paper is as follows: Section 2 provides a brief overview of the basics of the HI 21-cm signal around ionized bubbles. In section 3, we explain our methods to simulate the HI 21-cm signal around ionized bubbles, foreground contaminants, the generation of visibilities at baselines specific to a particular experiment, and system noise. Section 4 discusses foreground subtraction from the total visibilities and the application of the matched filter technique. Our results are presented in Section 5, while Section 6 presents a summary of the work undertaken in the paper alongside a discussion. Throughout our analysis we use the cosmological parameters $h=0.7$, $\Omega_{\rm m}=0.27$, $\Omega_{\Lambda}=0.73$, $\Omega_{\rm b}=0.044$, ${\sigma_{8}=0.83}$  and ${n_{\rm s}=0.96}$ consistent with the WMAP measurements [45].
@@ 3.1 HI 21-cm signal around ionized bubbles
Here, we briefly describe simulated ionized regions surrounding a very bright quasar at $z=7.1$, and around a collection of galaxies at $z=8.3$ during the EoR. First, we simulated the dark-matter distribution using the $N$-body code [^2] https://github.com/rajeshmondal18/N-body developed by [51]. We simulated this within a cubic box of the comoving size of $\sim 527$ Mpc with $2048^{3}$ grids and the same number of particles. The minimum halo mass resolved in this simulation is ${\sim 6.26\times 10^{9}M_{\odot}}$, corresponding to a minimum of $10$ dark matter particles, each with a mass of ${\sim 6.26\times 10^{8}M_{\odot}}$. We note that our simulations do not include the contributions of low-mass halos in the range $\bm{10^{8}M_{\odot}}$ to $\bm{\sim 6.26\times 10^{9}M_{\odot}}$. In reality, these halos are expected to play a significant role in the reionization of the IGM, potentially giving rise to more diffuse, complex, and smaller ionized structures during the EoR. Then we use a Friends-of-Friends  [^3] https://github.com/rajeshmondal18/FoF-Halo-finder (FoF) [52] halo finder to identify the collapsed halos that host the galaxies/bright quasars. These galaxies/quasars serve as the sources of ionizing photons during the EoR. Subsequently, the resolution of simulated cubes is degraded to $512^{3}$ grids. This resolution resembles the corresponding angular resolution of the uGMRT and SKA1-Low and has the added benefit of reduced computational requirements for simulation. We then employ a semi-numerical prescription [53, 54, 55] on these outputs to simulate ionization maps and the HI 21-cm differential brightness temperature maps.
@@
In the above two cases, the targeted HII bubbles are isolated and nearly spherical. This geometry arises because the ionizing radiation from other sources is negligible, and the IGM outside the targeted bubbles remains highly neutral. While these scenarios may be suitable at the onset of reionization, they are likely unrealistic for the later stages of the process. At later stages of reionization, ionized bubbles around very bright sources are surrounded by neighboring ionized regions, leading to an aspherical shape for the targeted bubble. The third scenario accounts for this complexity. The bottom panel of Fig. 1 displays a 2D map of the differential brightness temperature (${\delta T_{b}}$) at redshift $z=7.1$, where the central targeted HII bubble appears highly aspherical. The quasar at the center is the same as the first case. Surrounding it there are numerous other HII regions, unlike the first case. This results in a much lower mean neutral fraction, which is $0.52$ in this scenario.
@@
Figure 1: The top left and bottom panels show differential brightness temperature maps of the HI 21-cm signal around an ionized bubble of comoving size $23.5$ Mpc at $z=7.1$ corresponding to observing frequency of $175$ MHz. The top right panel shows a similar map around an ionized bubble of comoving size $27.8$ Mpc at $z=8.3$ corresponding to observing frequency of $153$ MHz. The mass-averaged neutral hydrogen fraction resulting from the galaxies outside the central HII region is $0.88$, $0.94$, and, $0.52$ for the top-left, top-right, and bottom panels respectively. The angular size of simulated maps is $\sim 3.3^{\circ}\times 3.3^{\circ}$ with $\sim 23^{\prime\prime}$ resolution.
@@ 3.2.1 Extragalactic radio point sources
One of the most significant foreground contaminants in HI 21-cm signal observation arises from the extragalactic radio point sources, such as Active Galactic Nuclei (AGN), quasars, and radio galaxies. At lower frequencies (${\sim 150}$ MHz), these sources dominate the sky at smaller angular scales $\lesssim 4^{\circ}$ [30]. Therefore, in order to study the HII regions at frequencies of $150$ and $175$ MHz, it is essential to model the point sources efficiently.
In our simulations, we model the population of extragalactic radio point sources using the following $150$ MHz differential source count model from [58]:
${\frac{dN}{dS}=\frac{10^{3.75}}{\rm Jy.Sr}\left(\frac{S}{1\,\rm Jy}\right)^{-1.6}},$ (3.3)
where $S$ is the flux density of the point sources. We consider the point sources to be in the flux range between $0.1$ mJy to $1$ Jy and they are randomly distributed across an angular size of ${\sim 3.3^{\circ}\times 3.3^{\circ}}$. Based on these parameters, we obtained $7775$ point sources from the simulation. We model the spectral energy distribution of the point sources using a simple power law (${S\propto\nu^{-\alpha}}$), where the spectral index $\alpha$ for each source is randomly assigned within the range of $0.7$ to $0.8$. The right panel of Figure 2 shows a 2D map of the point sources randomly distributed across an angular area of approximately $3.3^{\circ}\times 3.3^{\circ}$, with an angular resolution of ${\sim 23^{\prime\prime}}$.
@@ 3.4 Simulating visibilities
We generate the mock visibility data of the HI 21-cm signal and the foregrounds (DGSE and extragalactic point sources) at all simulated baselines at $256$ frequency channels, centered around frequencies $\nu=175$ MHz and $153$ MHz, using the Discrete Fourier Transformation (DFT) of the total simulated maps. For simplicity, we keep the baseline distribution the same in all frequency channels.
@@ 3.5 System noise
Our mock simulation considers $8$ hrs of observations with $16$s and $320$s integration time for the uGMRT and SKA1-Low respectively. This produces a total of $783000$ and $11773440$ baselines for the uGMRT and SKA1-Low respectively. Lower integration time will produce a higher number of baselines for the SKA1-Low and hence will increase the requirement of computational resources which is beyond our reach. The system noise contribution for $8$ hrs of observations will be too high to detect the HI 21-cm signal around the individual ionized bubbles. Therefore, we present our results for $2048$ and $3072$ hrs of uGMRT observations at redshifts $7.1$ and $8.3$, respectively, assuming a total of $256$ and $384$ nights of observations, with each night lasting $8$ hours. Further, for the third scenario, we present the results at redshift $7.1$ for $5120$ hours, assuming $640$ nights of observations. The increased observing time for the third scenario is due to the lower mean neutral hydrogen fraction outside the targeted bubble. Assuming all nights are the same, we reduce the system noise corresponding to an $8$ hrs observation by a factor of $16$, $19.6$, and $25.29$ for the three scenarios, respectively. The resulting mock signal becomes equivalent to $2048$, $3072$, and $5120$ hours of uGMRT observations. Similarly, for the SKA1-Low we present our results for $96$ hrs of observation for the first two cases at redshift $7.1$ and $8.3$ (neutral hydrogen fraction ${\sim 0.9}$) and for $200$ hrs of observation for the third case at redshift $7.1$ (neutral hydrogen fraction ${\sim 0.5}$). Eventually, we obtain mock visibility data with all the signal contributions (HI 21-cm signal, galactic foreground, extragalactic point sources, and system noise contribution) typical of the uGMRT and SKA1-Low observations.
@@
Figure 6 shows the total mock visibility comprising of HI signal, diffused galactic synchrotron emission, radiation from extra-galactic radio point sources, and system noise as a function of frequency for two baselines ${U=34\lambda}$ (left panel) and ${U=230\lambda}$ (right panel) for the uGMRT at $175$ MHz. We note that Figure 6 shows contribution both from the extragalactic point sources, galactic diffuse synchrotron radiation along with other components unlike Fig. 5 which shows galactic diffuse synchrotron emission only as the foreground component. The radio point sources are modeled using a simplistic power-law form as discussed in subsection 3.2.1. The green lines indicate visibilities due to point sources. We see that at a small baseline (left panel) the diffuse galactic synchrotron emission dominates over the point sources; at a large baseline (right panel), the point sources dominate over the diffuse galactic synchrotron emission. The total mock visibilities are similar for the second scenario at ${z=8.3}$ for both experiments, except for the fact that the system noise and foreground contributions are higher. Therefore, we do not show them here explicitly. The subsequent section discusses the subtraction of the foreground contribution.
@@
Figure 6: Left panel shows the total mock visibilities (HI 21-cm signal + DGSE + radio point sources + system noise) as a function of observing frequency for the uGMRT at baseline ${U=34\lambda}$ and the right panel shows the same at baseline ${U=230\lambda}$ at $175$ MHz observing frequency. Red dashed lines show diffuse galactic foreground contamination and the green lines show the point source contribution separately.
@@ 4.1 Foreground subtraction
Figure 7 presents the residual visibility at 175 MHz after the foreground contribution is subtracted from the total mock visibility for the uGMRT (top panels) and the SKA1-Low (bottom panels). Here, we have only considered the diffuse galactic synchrotron emission as the foreground contamination. However, we get a very similar residual signal when we perform similar analysis on the total mock visibilities considering the radio point source contribution, which has been shown in Figure 9. The residual signal contains the HI 21-cm signal, the system noise, and the un-subtracted foreground, if any. To assess the reliability of foreground subtraction, we compare the combined input HI 21-cm signal and system noise with the extracted signal for two different baselines ${U=34\lambda}$ (top left panel) and ${230\lambda}$ (top right panel), as depicted in Figure 8. In the bottom panels of Figure 8, we show a similar comparison for SKA1-Low for two different baselines ${U=11\lambda}$ (bottom left panel) and ${U=254\lambda}$ (bottom right panel). The plots show that the residual visibility closely matches the input visibility data, with only minor deviations. These slight deviations are possibly due to some residual foreground contribution and subtraction of the HI 21-cm signal and the system noise during the foreground subtraction. Consequently, we can conclude that we have mostly subtracted the foreground data from the mock visibility data, making it ready for further analysis, such as matched filtering. We do not present similar results for the second scenario at $z=8.3$ as it is similar to the first scenario.
Figure 8: These plots compare the recovered visibility (red dashed lines) after subtracting the diffused galactic foreground contribution with the original visibility consisting of the HI 21-cm signal and noise contributions (blue lines) for the uGMRT (top panels) and SKA1-Low (bottom panels) at same baselines and observing frequency considered in Figure 5. Here, the foreground component consists of diffuse galactic radiation only and visibilities are plotted for every fourth frequency channel for visual clarity.
Figure 9: These plots compare the recovered visibility (red dashed lines) after subtracting the extra galactic point sources with the original visibility consisting of the HI 21-cm signal and noise contributions (blue lines) for the uGMRT at same baselines and observing frequency considered in the top panels of Figure 5. Here, the foreground component consists of extra galactic point sources only. Visibilities are plotted for every fourth frequency channel for visual clarity.
@@ 5 Results
Here, we present our findings for the third case, where the targeted ionized bubble is surrounded by other ionized regions which result in a lower neutral hydrogen fraction of ${x_{\rm HI}=0.52}$. Further, the targeted bubble at the center also became aspherical. A lower neutral hydrogen fraction means that the background HI 21-cm signal is weaker, which results in a lower SNR value. We use the same matched filtering technique as described earlier but increased the observation time to $5120$ hours for the uGMRT and $200$ hours for the SKA1-Low, considering the lower SNR and increased difficulty in detecting ionized regions at a lower neutral fraction. Figure 16 shows the SNR as a function of filter size ${R_{\rm f}}$ for $5120$ hours of uGMRT observation (left panel) and $200$ hours of SKA1-Low observations (right panel) at $z=7.1$. As in previous cases, each curve represents a different noise realization. We see that for the uGMRT, the SNR peaks at a filter size ${\sim 25}$ Mpc, and for the SKA1-Low, it peaks at ${\sim 23}$ Mpc, with minor deviations in different noise realizations.
@@
Figure 16: This figure shows the signal-to-noise ratio for $10$ independent noise realizations corresponding to $5120$ hours of observation time for the uGMRT (left panel) and $200$ hours of SKA1-Low observations (right panel) at redshift $7.1$.
@@
Further, we see that peak SNR values in all ten realizations are ${\gtrsim 3}$ for the uGMRT and ${\sim 11}$ for the SKA1-Low. Like the previous cases, we assume a total bandwidth of $16$ MHz and $6.25$ MHz while subtracting the foreground for the uGMRT and SKA1-Low, respectively. We see that it is possible to reliably detect individual ionized bubbles in the HI 21-cm maps using the uGMRT with $5120$ hours of observation time and the SKA1-Low with only $\sim 200$ hours of observations at $z=7.1$ with significantly much lower neutral hydrogen fraction value.
@@
Figure 17: The left panel shows the mean signal-to-noise (SNR) ratio of $100$ independent noise realizations as a function of filter size for $5120$ hours of the uGMRT observations at redshift $7.1$. The right panel shows the SNR of $20$ independent noise realizations for $200$ hours of the SKA1-Low observations. The shaded region shows ${1\sigma}$ uncertainty estimated from independent realizations.
Figure 18: Same as Figure 12 at redshift $z=7.1$ with a lower neutral hydrogen fraction of $0.52$.
@@
The left panel of Figure 17 shows the mean SNR along with the ${1-\sigma}$ spread for $100$ noise realizations of the uGMRT, while the right panel shows the same for the SKA1-Low with $20$ realizations at $z=7.1$ for ${x_{\rm HI}=0.52}$. Further, the peak location varies between $16$ and $32$ Mpc for the uGMRT for different realizations (Figure 16), which can be seen in Figure 18, that shows a histogram of peak locations for $100$ realizations for the uGMRT. The peak value of the SNR for the uGMRT also changes in the range of $\sim 3-5$ and $10.5-12.5$ for the SKA1-Low.